solve this system of equations using either addition or substitution. thanks your help!!!!!!!!!
2m = n - 3
3m = 2n - 9
please show me how to check also!i keep messing up on that part also.

Answer by algebra help tutor(6)

Let's solve this system by using the process of elimination by subtraction. We will eliminate the n variable by multiplying the top equation by 2 and then subtracting the two equations.
2(2m=n-3)
3m=2n-9
4m=2n-6
- 3m=2n-9
m=3 Remember, -6-(-9) is -6+9, so the answer is 3.
To solve for n, substitute m=3 into any of the original equations.
2m=n-3
2(3)=n-3
6 =n-3
+3 +3 Addition property of equality
9 =n
So the answer is (3,9)
To check your answer, plug the point into both of the equations and see if they make sense.
2m=n-3 2(3)=9-3
6=6 checks out:)
3m=2n-9 3(3)=2(9)-9
9=18-9
9=9 checks out also:)